# Non-Linear Model-Based Sequential Decision-Making in Agriculture

### v1

@@ 5.3 Misspecified setting
truncated at $x_{0}=180$ and takes the same value $a+bx_{0}+cx_{0}^{2}$ for $x>x_{0}$. All other ingredients are identical to the well-specified experiment except that we increase the time horizon to $T=100$, specifically, price grid $p_{x}\in\{0.3,0.5,0.7\}$ \$ lb-1 N, yield price $p_{y}=\$5.00$ bu-1, action set $\mathcal{X}=\{0,50,100,\ldots,250\}$ lb N ac-1, noise $\sigma=0.5$, and the 5 algorithms: $\epsilon$-greedy, nonlinear-UCB, ViOlin, LinUCB, kNN-UCB. Ten independent replicates were run for each algorithm and price.
Figure 4: Comparing different Bandit algorithms for cumulative regret over time in the misspecified setting.
@@
Figure 4 plots the average cumulative regret when the fitted quadratic-plateau and Michaelis Menten model is misspecified with respect to the true Mitscherlich process. Relative to the well-specified case, the model-based regret curves shift upwards, reflecting the price of fitting an approximate response surface. Yet their slopes remain moderate, indicating that the fitted quadratic-plateau still guides the search toward profitable regions quickly. Note, that similar to the well-specified setting, the model-based non-linear bandits perform better than the linear and non-parametric baselines. Although the quadratic-plateau form cannot capture the exponential approach to an asymptote exactly, its concave-then-flat shape matches the broad geometry of the Mitscherlich curve. As a consequence, all three nonlinear model-based algorithms, $\epsilon$-greedy, UCB, and ViOlin continue to accumulate substantially less regret than the linear (LinUCB) or non-parametric (kNN-UCB) baselines. This shape-compatible misspecification warrantees a more structured theoretical and experimental research, but our preliminary results are promising in this direction. Although ViOlin’s greedier, curvature-penalized score again avoids the large initial dip seen in UCB and $\epsilon$-greedy, note that UCB and $\epsilon$-greedy seem to perform comparably, especially in the Michaelis Menten fit, perhaps suggesting that exploration might help in misspecified settings. Other details can be found in the Appendix.
@@ 6 Conclusion and future work
In this work, we developed and systematically evaluated model-based nonlinear bandit algorithms for optimizing input decisions in agricultural systems. By framing the nitrogen rate optimization task as a sequential decision problem, we demonstrated that mechanistic model-based approaches, such as those relying on Mitscherlich, quadratic plateau, Michaelis-Menten, or logistic dose-response curves, can be flexibly integrated into the bandit framework to yield interpretable and effective decision-making strategies.
Our extensive simulation study revealed two principal findings. First, when the fitted model matches the true data-generating process, model-based bandit algorithms achieve substantially lower regret than nonparametric and linear baselines, highlighting the value of mechanistic knowledge. Second, when the fitted model is misspecified relative to the true reward structure, we observed a notable increase in regret, with the degree of performance loss depending on the extent of model mismatch. Notably, even under moderate misspecification, model-based bandits can still outperform linear methods such as LinUCB or fully model-free methods such as kNN-UCB, particularly when the mechanistic model provides a reasonable approximation of the underlying process and there are limited samples to learn from.
A key theoretical contribution of this work is the establishment of a conservative sample complexity bound for the ViOlin algorithm, which guarantees that the sample complexity, i.e., the number of rounds required to learn the local maxima of a non-linear reward function, scales polynomially in the desired accuracy and is independent of the dimension of the arm space. Although this result could be conservative for simpler models like we have considered, such applications further appeal to developing theoretical and methodological research that focuses on sample efficiency along with regret minimization. This is especially important in real-world applications like agriculture, where decision-makers often face tight resource constraints and require efficient learning algorithms that make the most out of limited data.
Our results underscore the importance of model selection and the trade-off between interpretability and robustness to misspecification in real-world sequential decision-making. Furthermore, we emphasize the indispensable role of domain experts in guiding data-driven algorithms, both by providing informed priors and initial conditions and by aiding in the interpretation of estimated parameters. Such collaborations ensure that adaptive management strategies remain both scientifically grounded and practically relevant.
Looking forward, several promising directions remain for future research. First, a systematic investigation into model selection or model averaging within the bandit setting could mitigate the risk of misspecification and enhance robustness. Extending the current framework to multi-dimensional or spatially varying decisions (i.e.,varying nitrogen rates across heterogeneous fields) represents an important direction, particularly given the prevalence of spatial heterogeneity in agricultural systems. Another practical extension is to incorporate weather and soil conditions as covariates in the non-linear models as these can play a significant role in the management decisions being made. Related to this is the fact that the non-linear models could be non-stationary and evolve over time. Developing strategies that adapt to non-stationarity in small sample settings is another relevant direction for future work. Another compelling direction is to bridge mechanistic and machine learning approaches, such as integrating semi-parametric or hybrid models that combine the interpretability of domain knowledge with the flexibility of data-driven components. Further, establishing tighter theoretical guarantees for regret under shape-constrained model misspecification, and extending these results to settings with delayed, censored, or partially observed rewards would further strengthen the practical applicability of bandit algorithms in complex, real-world domains.
From a broader perspective, the development of sample-efficient, interpretable, and robust sequential decision-making algorithms is crucial for advancing precision agriculture and achieving the Sustainable Development Goals (SDGs). These methods are particularly impactful for smallholder farmers operating with limited data, as they enable principled data-driven decision-making even under uncertainty and resource constraints. Overall, our study lays the groundwork for principled, adaptive management strategies in agriculture and beyond, with the potential for significant societal and environmental impact.
Disclosure Statement
@@ Appendix D Additional Simulation Results
In this section, we present the simulation results for other non-linear models that we considered in the simulation experiment of Section 5. First for the well-specified setting, we present the results for the Logistic and Mitscherlich model that was fit instead of Quadratic Plateau model. We consider the following true parameters for the simulation settings for each of the three models:
  • Logistic: $A=120,B=0.05,C=125,d=70$
  • Mitscherlich: $A=120,b=0.015,d=80$

### v2

@@ 5.3 Misspecified setting
truncated at $x_{0}=180$ and takes the same value $a+bx_{0}+cx_{0}^{2}$ for $x>x_{0}$. All other ingredients are identical to the well-specified experiment except that we increase the time horizon to $T=100$, specifically, price grid $p_{x}\in\{0.3,0.5,0.7\}$ \$ lb-1 N, yield price $p_{y}=\$5.00$ bu-1, action set $\mathcal{X}=\{0,50,100,\ldots,250\}$ lb N ac-1, noise $\sigma=0.5$, and the 5 algorithms: $\epsilon$-greedy, nonlinear-UCB, ViOlin, LinUCB, kNN-UCB. 10 independent replicates were run for each algorithm and price.
Figure 5: Average cumulative regret trajectories across 10 replications in the misspecified setting. Nonlinear model-based algorithms still consistently outperform linear and nonparametric baselines.
@@
Figure 5 plots the average cumulative regret when the fitted quadratic-plateau and Michaelis Menten model is misspecified with respect to the true Mitscherlich process. Relative to the well-specified case, the model-based regret curves shift upwards, reflecting the price of fitting an approximate response surface. Yet their slopes remain moderate, indicating that the fitted quadratic-plateau still guides the search toward profitable regions quickly. Note, that similar to the well-specified setting, the model-based non-linear bandits perform better than the linear and non-parametric baselines. Although the quadratic-plateau form cannot capture the exponential approach to an asymptote exactly, its concave-then-flat shape matches the broad geometry of the Mitscherlich curve. As a consequence, all three nonlinear model-based algorithms, $\epsilon$-greedy, UCB, and ViOlin continue to accumulate substantially less regret than the linear (LinUCB) or non-parametric (kNN-UCB) baselines. These results highlight the resilience of shape-compatible models under misspecification and point to rich opportunities for deeper theoretical and empirical study. Although ViOlin’s greedier, curvature-penalized score again avoids the large initial dip seen in UCB and $\epsilon$-greedy, note that UCB and $\epsilon$-greedy seem to perform comparably, especially in the Michaelis Menten fit, perhaps suggesting that exploration might help in misspecified settings. Other details can be found in the Appendix.
@@ 6 Conclusion and discussion
In this work, we developed and systematically evaluated model-based nonlinear bandit algorithms for optimizing input decisions in agricultural systems. By casting the nitrogen rate optimization task as a sequential decision problem, we demonstrated that mechanistic model-based approaches, such as those relying on Mitscherlich, quadratic plateau, Michaelis-Menten, or logistic dose-response curves, can be embedded into adaptive learning algorithms. This integration yields decision rules that are not only sample-efficient but also interpretable, biologically grounded, and directly actionable for practice.
Our simulation studies revealed two central insights. First, under well-specified conditions, model-based bandits achieve substantially lower regret than linear or nonparametric benchmarks, underscoring the advantage of exploiting mechanistic structure. Second, under model misspecification, regret inevitably increases, but mechanistic models still provide competitive performance when the assumed shape is qualitatively correct, outperforming generic baselines in small-sample regimes. This robustness is critical in agriculture, where each additional trial represents both a financial cost and an environmental burden.
A key theoretical contribution of this work is the establishment of dimension-independent sample complexity guarantees for the ViOlin algorithm. Because each field trial corresponds to real costs of time, labor, and inputs, the ability to guarantee learning under strict resource constraints is as valuable as the specific rate of convergence. We show that the number of rounds required to identify a near-optimal fertilizer rate scales polynomially with the desired accuracy and is independent of the size of the action space. These guarantees provide a principled baseline for nonlinear model-based bandits, ensuring that learning remains sample-efficient even in high-dimensional decision settings. These bounds illustrate the theoretical tractability of model-based approaches for structured agronomic models and establish a foundation for sharper, model-specific analyses. Importantly, our simulations suggest that much faster convergence is possible in practice, underscoring a fruitful gap between theory and application that motivates further research.
Our findings highlight model choice and robustness as equally critical to algorithmic performance. The trade-off between interpretability and flexibility is especially acute in real-world sequential decision-making: mechanistic models provide transparency and scientific grounding, while more flexible models offer resilience to misspecification. We emphasize that collaboration with domain experts is indispensable for resolving this tension. Agronomic knowledge can inform priors, parameter ranges, and plausible initial values, while also guiding the interpretation of fitted parameters. Such integration ensures that adaptive management strategies are not only statistically principled but also credible and actionable for practitioners.
Looking forward, several avenues offer promising directions for future research. While our results are based on controlled simulations, future work will involve applying these methods to field experiments and real-time on-farm decision making. Such studies are essential for validating the algorithms under practical constraints, including delayed feedback, heterogeneity across fields, and limited opportunities for intervention. Developing principled strategies for model selection or model averaging within the bandit setting could mitigate misspecification risks. Extending the framework to multi-dimensional or spatially varying treatments (e.g., variable-rate nitrogen across heterogeneous fields) would better reflect on-farm conditions. Incorporating weather, soil, and other covariates into nonlinear reward models, and designing algorithms that adapt to non-stationary environments, are further priorities. Methodologically, bridging mechanistic and machine-learning approaches through hybrid or semi-parametric models could combine interpretability with flexibility. From a theoretical perspective, refining regret bounds for structured nonlinear models and extending them to settings with delayed or censored rewards would strengthen the practical utility of these algorithms.
From a broader perspective, this work contributes to the development of sample-efficient, interpretable, and robust sequential decision-making methods that can directly inform precision agriculture. By enabling smallholder farmers and resource-constrained producers to learn from limited experiments, such methods support progress toward Sustainable Development Goals, in particular SDG 2 (Zero Hunger) and SDG 12 (Responsible Consumption and Production). Overall, our study lays the groundwork for adaptive experimentation strategies that advance both the methodological frontiers of sequential decision-making while contributing to the societal goals of sustainable, data-efficient agriculture.
Reproducibility statement
All simulations were implemented in Python 3.11 using standard scientific libraries (NumPy, SciPy, Matplotlib). We report the data-generating processes, parameter values, and hyperparameters in full within the paper and Appendix, including initial values used for nonlinear least-squares estimation, fertilizer and grain price settings, and noise levels. Each experiment was repeated independently across 10 replicates with fixed randomization schemes to assess variability. Our simulation code (anonymized for review) will be made publicly available upon publication to facilitate exact replication of results and adaptation to new scenarios.
Data availability statement
The simulation code and data used in this study will be made publicly available on Github upon publication.
Ethics statement
This study does not involve human participants, human data, or animals, and therefore ethics approval was not required.
Funding statement
This research received no specific grant from any funding agency in the public, commercial, or not-for-profit sectors.
Conflict of interest
@@ Appendix C Baseline Algorithm descriptions
C.1 Toy Illustration of Algorithm Behavior
To make the exploration–exploitation logic more tangible, we provide a simple illustration using a Mitscherlich yield function
$Y(x)\;=\;80+120\,\big{(}1-e^{-0.015x}\big{)},$
with crop price $p_{y}=5$ per unit yield (bu-1), fertilizer cost $p_{x}=0.7$ per unit (lb-1 N), and action grid $\{0,50,100,150,200\}$ lb N ac-1. Table 3 shows four rounds of decision making under three algorithms: $\epsilon$-Greedy, UCB, and ViOlin.
Table 3: Toy demonstration of algorithm behavior (profit in \$).
Round $\varepsilon$-Greedy UCB ViOlin
Action Profit Action Profit Action Profit
1 Explore 50 541 Explore 100 610 Explore 50 540
2 Explore 150 650 Explore 150 648 Exploit 150 649
3 Exploit 150 648 Exploit 150 647 Exploit 150 648
4 Explore 50 541 UCB $\to$ 200 645 Exploit 150 647
This toy comparison highlights the distinct decision logics:
  • $\epsilon$-Greedy alternates between random exploration and exploiting the currently estimated best dose.
  • UCB incorporates an explicit uncertainty bonus, occasionally sampling higher but uncertain doses (round 4).
  • ViOlin emphasizes curvature and stability, quickly converging to the plateau dose and avoiding unnecessary exploration.
Even in this simple 4-round setting, the three strategies exhibit clearly different exploration styles. Such illustrations help convey to applied readers that algorithm choice affects not only long-run performance, but also the sequence of recommendations farmers may observe in practice.
@@ Appendix D Additional Simulation Results
In this section, we present the simulation results for other non-linear models that we considered in the simulation experiment of Section 5. First for the well-specified setting, we present the results for the Logistic and Mitscherlich model that was fit instead of quadratic plateau model. For reproducibility, Table 4 reports the true parameter values used to generate data in our simulations, alongside the initial values supplied to the nonlinear least-squares estimation routines. These initialization choices were selected to be reasonably close to the true parameters, reflecting what might be obtained from domain knowledge or prior agronomic experiments. As discussed in Section 5, good initialization improves convergence in small-sample regimes, though all algorithms eventually stabilize near the true values as data accumulate.
Table 4: True and initial parameter values used in simulation studies.
Model True Parameters Initial Values
Mitscherlich A=120,b=0.015,d=80A=120,\ b=0.015,\ d=80 A=100,b=0.01,d=75A=100,\ b=0.01,\ d=75
Quadratic Plateau $a=80,\ b=1.2,\ c=-0.003,\ x_{0}=180$ $a=75,\ b=1.0,\ c=-0.002,\ x_{0}=160$
Michaelis–Menten a=150,b=100,d=60a=150,\ b=100,\ d=60 a=120,b=80,d=50a=120,\ b=80,\ d=50
Logistic A=120,B=0.05,C=125,d=70A=120,\ B=0.05,\ C=125,\ d=70 A=100,B=0.03,C=100,d=65A=100,\ B=0.03,\ C=100,\ d=65